OpenStudy (anonymous):

What scale did you use to make your model? How did you decide to use this scale? Show your calculations. Based on the scale you used, how long (in centimeters) was the Precambrian period compared to the rest of the scale? How many “centimeters” of time separated the dinosaurs and humans on Earth? Were you surprised by the length of the time scale when you saw it on a continuous strip of paper? Why or why not? Explain your answer as it relates to the scope of Earth’s geologic past.

jimthompson5910 (jim_thompson5910):

this sounds like a continuation of a prev problem

jimthompson5910 (jim_thompson5910):

and there's info missing (it seems)

OpenStudy (anonymous):

It also sounds like something you're intended to do yourself. They're looking for your opinions, justifications, and reactions.

OpenStudy (anonymous):

1.What scale did you use to make your model? 2.How did you decide to use this scale? Show your calculations. 3.Based on the scale you used, how long (in centimeters) was the Precambrian period compared to the rest of the scale? 4.How many “centimeters” of time separated the dinosaurs and humans on Earth? 5.Were you surprised by the length of the time scale when you saw it on a continuous strip of paper? Why or why not? Explain your answer as it relates to the scope of Earth’s geologic past.

jimthompson5910 (jim_thompson5910):

There's missing info because they didn't provide when the precambrian period was (or how long it was)

OpenStudy (anonymous):

Cut a piece of adding-machine tape 4.6 meters long. (If you have only a ruler, remember that one meter equals 100 centimeters or 39.37 inches.) If no adding-machine tape is available, you can cut plain paper into long strips, each about 5 centimeters wide, and attach them to make a continuous piece of paper, approximately 4.6 meters long. Note: You may also create a digital time scale using a word-processing or spreadsheet program of your choice. Choose this option if you do not have access to a digital camera. Determine an appropriate scale for placing the geologic time scale of 4.6 billion years (4,600 million) on the piece of adding-machine tape. If you need help creating the scale, try the scale calculator. Plot each era and period on the adding machine tape using the scale you chose from your calculations. As you indicate each era or period, make notes on the paper about significant events that occurred during that time. Events can be biological or geological in nature.

jimthompson5910 (jim_thompson5910):

so which part isn't making sense

OpenStudy (anonymous):

all :(

jimthompson5910 (jim_thompson5910):

you basically have to divide all of geologic time into 4.6 meters So think of a timeline that is 4.6 meters long and each "tick" represents a specific event

jimthompson5910 (jim_thompson5910):

eg: the starting tick is the creation of the earth

OpenStudy (anonymous):

then after that what do i do :(

jimthompson5910 (jim_thompson5910):

or the middle tick is some period in the middle point of earth's age

jimthompson5910 (jim_thompson5910):

and the last tick would be modern day

jimthompson5910 (jim_thompson5910):

of course you would have other ticks as well

jimthompson5910 (jim_thompson5910):

where you place these events depends on how long ago each event happened (and the time space between each event)

OpenStudy (anonymous):

so whaT woUld the Answers be :)

OpenStudy (anonymous):

There aren't just answers, it's a project. You have to do the project. Yourself.

jimthompson5910 (jim_thompson5910):

there's a lot to this project than simple answers like "yes" or "23" and unfortunately I don't know much about geology to know all the answers to this long question

jimthompson5910 (jim_thompson5910):

yes and it's a project where you have to create a timeline, so that's not really an "answer"

OpenStudy (anonymous):

Which of the following equations represents a perpendicular line to 6x - 2y = 20?

OpenStudy (anonymous):

Jackie conducted a survey of her fellow students in the freshman class to determine what flavor of ice cream they prefer to be served at their end of the year party: vanilla, chocolate, or strawberry. She made the Venn diagram below to show her class sponsor. http://learn.flvs.net/webdav/assessment_images/educator_algebra1_v14/10_18_10_s1.gif How many students belong to the set ∼C ∪ (S ∩ ∼V)? 332 403 531 585

jimthompson5910 (jim_thompson5910):

Solve for y to get 6x - 2y = 20 -2y = -6x+20 y = 3x-10 The slope here is 3 So anything with a slope of -1/3 will be perpendicular